### v1

# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

Abstract
We present a minimal operator-only term rewriting system with seven constructors and eight reduction rules. Our main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. From strong normalization, we derive a certified normalizer with proven totality and soundness. Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment. We establish impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication. The work demonstrates fundamental limitations in termination proving for self-referential systems. We state a conjecture: no relational operator-only TRS can have its full-system termination proved by internally definable methods. Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting. Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures. This structural limitation applies to any TRS expressive enough to encode ordered data. All theorems have been formally verified in a proof assistant. The formal development is available to program committee members and referees upon request for purposes of peer review.
1 Introduction
We develop a minimal operator-only rewrite calculus (KO7): the object language contains only constructors and operators with rewrite rules; there are no binders, types, external axioms, or semantic predicates. The rules are the semantics. Our goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) (optionally) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture that no operator-only TRS capable of ordered computation can have its full-system termination proved by internally definable methods.
Scope. All formal results are established for a guarded safe subrelation. We do not claim a single global measure for the full unguarded relation. Moreover, we exhibit a precise negative: at the root peak $\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$, local join fails in the full relation. Therefore the full relation is not locally confluent at that peak, hence not confluent. The designed remedy is to work in the safe fragment with guarded/context joiners (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If we define an internal provability predicate by “$t$ reduces to $\top$”, then under SN the set $\{t\mid t\Rightarrow^{\ast}\top\}$ is decidable by normalization/backtracking. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; the right move is stratification.
Contributions.
(1) A duplication-robust SN proof for KO7 using a triple-lex measure with a multiset (DM) component and an MPO-style head precedence; (2) a total, proved-correct normalizer; (3) a guarded Newman module for the safe relation that yields confluence (hence unique normal forms) from SN + local confluence; (4) a decidability result for reachability under SN; (5) a catalog of impossibility results for additive and polynomial measures under duplication; (6) a formal verification of all results; (7) a conjecture on undecidable termination for operator-only systems within a fixed base theory.
Highlights (formalization summary).
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz–Manna multiset rank, and an ordinal.
  • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
  • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
  • Full Step caveat: We exhibit a specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails, justifying the restriction to the safe subrelation.
  • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
2 Background: TRSs, SN, reachability, and Newman
We assume standard abstract reduction and term rewriting notions [BaaderNipkow98, Terese03]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [Newman42], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [BaaderGiesl24]. Modularity holds in certain linear, non-collapsing combinations [CaronCoquide94], while even flat non-linear systems exhibit undecidable reachability and confluence [Mitsuhashi06]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [DershowitzManna79] or (recursive) path orders [Dershowitz87].
3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including an explicit case split on eqW). For concreteness we list the rule shapes below.
Rule Head Arity Shape Dup?
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 1: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing. R5 redistributes $s$ to two RHS positions.
Tiny example (trace consequences).
Using the verified normalizer, we observe:
  • Integrate/delta: $\mathrm{integrate}(\delta\,t)\Rightarrow^{\ast}\mathrm{void}$ (verified).
  • Equality (meta-level consequence under confluence):
    • If $\mathrm{nf}(a)=\mathrm{nf}(b)$ then $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{void}$.
    • Otherwise $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, confluence ensures these outcomes are unique.
Note. We prove SN and Newman-based confluence for the safe fragment.
4 Strong normalization
We define a triple-lexicographic measure
$\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz–Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are stated explicitly (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). In duplicating branches (if extended to broader systems) we would use a compact MPO-style head precedence.
Step vs SafeStep (measure unification).
For the SafeStep relation we use a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$. For the full relation we only provide a disjunctive decrease certificate: each step is covered by either a KO7 lex drop or an MPO triple drop. We do not claim a single global well-founded measure for all full steps.
Theorem 1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}(t)$.
[…]
DM vs MPO on rules (explicit).
For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$. For eqW, we use a compact MPO-leaning measure where the head precedence orients $\mathrm{eqW}\,a\,b$ strictly above $\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
  • Base-order premise (merge-cancel): in $\mathrm{merge}\,t\,t\to t$, the RHS is strictly smaller than the LHS.
  • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
No-Go for constant bumps on $\kappa$ (generic duplicator).
Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
Let $\kappa$ be a max-depth-style counter and fix any $k\in\mathbb{N}$. For a duplicating rule of the shape $r(S)\to C[S,S]$ (one redex replaced by two occurrences of a subterm $S$), there exists an instance where $\kappa(\text{LHS}){+}k=\kappa(\text{RHS}){+}k$, so no strict lex drop occurs. The same holds for a boolean phase flag alone.
Proof sketch. Choose $S$ with $\kappa(S)\geq 1$ and let $\mathrm{base}$ bound the context. Then $\kappa(\text{LHS})=\mathrm{base}{+}1$ while $\kappa(\text{RHS})=\max(\kappa(C[S]),\kappa(C[S]))=\mathrm{base}{+}1$. Adding a fixed $k$ preserves equality; a single flag does not alter the tie. $\square$
Remark 1.
[…]
Duplication stress identity.
For any additive counter $\rho$ that counts a single removed redex and sums subpieces, generic duplicators satisfy
$\rho(\text{after})=\rho(\text{before})-1+\rho(S),$
so there is no strict drop when $\rho(S)\geq 1$. The robust fix uses DM/MPO: replace one element by a multiset of strictly smaller elements (DM), or use RPO/MPO with a precedence/status such that the LHS redex strictly dominates each RHS piece.
Corollary 1 (Strong normalization).
The guarded relation SafeStep for KO7 is strongly normalizing.
Genealogy of failures (why DM/MPO).
We record minimal counterpatterns that motivate the multiset/path components:
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 1).
  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
  • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
The fix is DM/MPO: ensure each RHS piece is strictly smaller than the removed redex in a base order and lift via a multiset/path extension.
5 A certified normalizer
By well-founded recursion on $\mu^{3}$ we define a normalization function. We prove the following properties (separated to avoid overfull lines):
(Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
(Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
These are exposed in the formalization. We make no efficiency claim: worst-case normalization cost follows the termination witness in play (cf. the small-term reachability bounds in [BaaderGiesl24]). In confluent systems, decision reduces to one normalization and an equality check.
6 Local confluence and Newman (guarded safe relation)
We discharge local confluence by joining the finite set of critical pairs (when present). Combining Cor. 1 with Newman’s Lemma [Newman42] yields:
Theorem 2 (Confluence and unique normal forms).
If a relation is strongly normalizing and locally confluent, then it is confluent. Hence every term reduces to a unique normal form.
Instantiation (SafeStep). Combining Cor. 1 (SN for SafeStep) with local-join lemmas yields confluence and unique normal forms for the safe fragment by Newman’s Lemma. In contrast, the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ under $\kappa^{M}(a)=0$; thus full confluence does not hold. The SafeStep relation restricts eqW to cases where arguments are guarded, preventing this divergence.
The star-star join proof follows the standard accessibility (Acc) recursion at the source, with a case split over star shapes (“head step + tail”) and composition via transitivity. The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
Scope and Guarantees (formalization-accurate).
We work with a guarded safe subrelation SafeStep for which we prove:
  • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local Confluence (SafeStep). We provide local-join lemmas per root shape and context wrappers. Newman then yields confluence and unique NFs for the safe fragment.
  • Full Step per-rule decreases (Hybrid). For each kernel rule, there is a per-step decrease witnessed either by KO7’s $\delta/\kappa^{M}/\mu$ lex or by an MPO-leaning $\mu$-first triple. A uniform global aggregator for all of Step is left as future work.
Critical-pair coverage.
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which we show is not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta  (unique target void)
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left  (unique target $t$)
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right  (unique target $t$)
$\mathrm{merge}\;t\;t$ localJoin_merge_tt  (unique target $t$)
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero  (unique target $b$)
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ  (unique target $\mathrm{app}\,s\,(\mathrm{rec}\,\Delta\,b\,s\,n)$)
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne  (unique target $\mathrm{integrate}(\mathrm{merge}\,a\,b)$)
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ not locally joinable at root
Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
$\delta$-guard: definition and decidability.
Define the safe-phase predicate by $\delta\!\operatorname{-guard}(t)\iff\delta\!\operatorname{Flag}(t)=0$. Here $\delta\!\operatorname{Flag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\delta\!\operatorname{Flag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\delta\!\operatorname{Flag}(t)=0$.
[…]
$\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
7 Impossibility results
We formally establish the failure of several simpler measure strategies, which necessitate the use of DM multiset orders or MPO. These negative results are verified in the formal development.
  • Additive bumps fail: No measure of the form $\kappa(t)+k$ strictly decreases on generic duplicating rules.
  • Bare flags fail: A single boolean flag is insufficient to orient rules that lift depth.
  • Polynomial interpretations fail: Any polynomial interpretation involving fixed constants requires external arithmetic axioms to “orient” the rule, violating the operator-only constraint.
Polynomial Impossibility.
Consider polynomial interpretations of the form $M(t)\in\mathbb{N}$ for a duplicating rule $f(x)\to g(x,x)$. A polynomial measure assigning $M(f(x))=M(x)+p$ requires $M(x)+p>2M(x)$ for strict decrease, implying $p>M(x)$. This fails for unbounded $M(x)$.
More generally: any polynomial proof for a system with duplicating rules must hardcode specific constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the proof. This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
Internally definable measures.
We use a simple contract for internally defined termination measures to structure these negative results:
Definition 1 (Internally definable measure).
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
8 Decidability of Reachability
Theorem 3 (Fixed-target reachability).
In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
This connects our work to classical results: if we could encode undecidable properties through reduction to constants, we would violate known theoretical limits. This motivates stratified approaches in proof assistants, where object-level and meta-level reasoning are carefully separated.
Assumptions (model).
We work with a finite first-order TRS over finite terms. The decision method is: compute a normal form (by SN) and check whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
Complexity context.
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [BaaderGiesl24]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [CaronCoquide94], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [Mitsuhashi06].
9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means.
A relational operator-only TRS is, by definition, one capable of ordered computation. The terms are equivalent: “relational” = “capable of ordered computation.” Such systems can express iteration over successors, comparison of structure depths, or any form of sequential counting. The minimal signature for such capability includes a recursor, an operator that applies a step function iteratively across a successor-structured argument. Examples include primitive recursion over natural numbers, fold over lists, and any construct that steps through an ordered index. A TRS that cannot express such iteration (e.g., a flat pattern-matching system with no recursive structure) is not relational in this sense.
The critical property of relational systems is that their recursor redistributes its step argument across recursive calls. This redistribution is not optional: it is what makes ordered computation possible. But it also defeats all additive termination measures.
The structural barrier.
Any operator system capable of ordered computation requires a recursor. A recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 2 (Internally definable measure).
An internally definable termination method for signature $\Sigma$ uses only: simplification orders with fixed precedence/status on $\Sigma$ (LPO/RPO/MPO), DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Conjecture 1 (Full Termination Conjecture for Relational TRSs).
No relational operator-only TRS can have its full-system termination proved by internally definable methods. Specifically: let $R$ be an operator-only TRS with a recursor rule of the form
$\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n)).$
No internally definable measure (Definition 2) proves termination of $R$ when the step argument $s$ is unrestricted. The structural redistribution of $s$ into both the function application and the recursive call defeats every additive, polynomial, or path-order measure that does not import external arithmetic or ordinal axioms. Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
Scope of the conjecture.
The conjecture applies to any TRS expressive enough to encode ordered data (natural numbers, lists, trees with successor structure, or any type supporting iteration). Systems without such structure (purely ground, non-recursive, or lacking a step-iteration pattern) fall outside the conjecture’s scope. The claim is that once a system has enough structure for ordered computation, its full termination escapes internal proof methods.
Evidence.
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta$-phase bit restricts allowable step arguments; without such restriction, the full system admits unbounded recursive depth. The gap between internally definable ranking functions and what would be required cannot be closed without external methods (axioms, encodings, or borrowed arithmetic).
10 Formalization structure
The formal verification is implemented in a proof assistant. The project structure separates the kernel definitions, meta-theory proofs, and impossibility results:
  • Termination proofs: Establishes the per-rule decreases and strong normalization for the safe fragment using the triple-lexicographic measure.
  • Normalization: Defines the normalization function and proves its totality and soundness.
  • Confluence: Implements the Newman engine and the star-star join, relying on local join lemmas.
  • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available to reviewers under appropriate confidentiality agreements.
Appendix: Addendum
KO7 Rules Table (7 constructors, 8 rules)
Rule Head Arity Shape Dup?
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 2: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing (erases one copy). R5 redistributes $s$ to two positions on the RHS; see Remark below.
Triple-lexicographic measure and duplication handling (DM/MPO)
We order $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ by lex:
  • $\delta$-flag: phase bit dropping on the successor recursion branch.
  • Multiset $\kappa^{M}$: a Dershowitz–Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
  • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
Aggregation for $\kappa^{M}$ (union, not sum).
We emphasize that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel we use a $\kappa$-branch: when $\kappa^{M}(a)\neq 0$ we obtain a left-lex drop via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
Short witness snippets (toy duplication).
Toy rule: $\mathrm{pair}(\mathrm{s}\,x,\,y)\to\mathrm{pair}(x,\,\mathrm{pair}(y,\,y))$.
DM multiset on sizes. Let $S(x,y)=\mathrm{size}(\mathrm{pair}(\mathrm{s}\,x,\,y))=\mathrm{size}(x)+\mathrm{size}(y)+2$. Then $\mathrm{size}(x)<S(x,y)$ and $\mathrm{size}(y)<S(x,y)$. Use $X=\varnothing$, $Y=\{\mathrm{size}(x),\mathrm{size}(y)\}$, $Z=\{S(x,y)\}$ to conclude $Y<_{\mathrm{DM}}Z$.
MPO triple weight. $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
[…]
Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star–star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly we do not claim full-step confluence.
Module map
The verification logic is distributed across key modules for SN, normalizer, and confluence engine. Local confluence lemmas are proved separately. Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
[…]
Rule Base component Precedence/Status Witness Source
merge void left $\mu$ (ordinal) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) Theorem 4.2 Termination
merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
rec succ $\delta$ phase bit (1 $\to$ 0) Theorem 4.5 Termination
eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
integrate (delta t) $\mu$ (unique target) Theorem 4.8 Termination
toy duplication DM on size s >> pair Lemma 7.1 Impossibility
toy duplication MPO triple headRank(s) >> headRank(pair) Lemma 7.2 Impossibility
Table 3: Per-rule orientation summary.
Parenthetical aliases: for substitution convenience we expose simp forms alongside the main rec lemmas.
Local-join report (critical peaks and unique-target cases)
Shape Join lemma name Notes
merge void void localJoin_merge_void_void both branches to void
merge t t localJoin_merge_tt joins at $t$ (context-lifted)
integrate (delta t) localJoin_int_delta unique target void
[…]
eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies (unique target)
eqW a a, guard kappaM a $\neq$ 0 localJoin_eqW_refl_guard_ne refl blocked; only diff
merge a b (no void, a $\neq$ b) localJoin_merge_no_void_neq no root step; vacuous
integrate non-delta localJoin_integrate_non_delta no root step; vacuous
[…]
Table 4: Local-join witnesses for the safe relation (root).
Ctx-local join wrappers
When needed in context, the root local-join lemmas lift to context via wrappers:
Shape Ctx wrapper lemma Notes
merge a b (no void, a $\neq$ b) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
eqW a a, kappaM guard localJoin_ctx_eqW_refl_guard_ne refl blocked
[…]
eqW (delta n) (delta n) localJoin_ctx_eqW_refl_when_a_is_delta literal delta case
Table 5: Ctx-local join wrappers.
Appendix: Orientation Strategy
This summarizes the exact orientation strategy realized in the proofs (no global RPO required):
  • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$, wrapped by drop_R_rec_zero.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0) via drop_R_rec_succ.
  • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.
  • Conceptual precedence (if one prefers an MPO reading) aligns with rec >> merge >> app >> integrate >> eqW >> delta >> void, but the formal proofs do not require a global precedence declaration.
Formal anchors:
  • Merge void left/right: $\mu$-first or DM tie plus $\mu$.
  • Merge cancel: DM multiset drop on $\kappa^{M}$.
  • Rec zero: drop_R_rec_zero with dm_drop_R_rec_zero for the inner DM piece.
  • Rec succ: drop_R_rec_succ via the $\delta$ phase bit (1 $\to$ 0).
  • EqW refl/diff: $\mu$-first MPO triple.
  • Integrate delta: unique target void.
Appendix A Related work
We follow standard expositions in Baader-Nipkow and Terese [BaaderNipkow98, Terese03]. Termination under duplication relies on DM multiset orders [DershowitzManna79] and (recursive) path orders [Dershowitz87]. Newman’s Lemma is used in its classic ARS form [Newman42]. For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [Goodstein44, KirbyParis82, Buchholz86]. For fixed-target reachability complexity and modularity results, see [BaaderGiesl24, CaronCoquide94]; for undecidability at shallow depth in non-linear systems, see [Mitsuhashi06].
Appendix B Conclusion
We gave an operator-only kernel and proved strong normalization and confluence for the guarded safe subrelation, yielding a certified normalizer and unique normal forms via Newman’s Lemma. We also established an explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$), blocking confluence for the full relation; this motivates the designed SafeStep fragment. A decidability result explains why single-level Gödel encodings clash with global termination. We proved a simple impossibility for additive bumps under duplication and certified DM/MPO orientations. Future work: complete a stratified meta-layer; develop arithmetic encodings within KO7; and map the exact boundary between internally definable measures and external ordinal strength.