### v3

# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

Abstract
This paper presents a minimal operator-only term rewriting system (KO7) designed to investigate the foundations of self-verifying computation. The system consists of seven constructors and eight reduction rules, with no external axioms, no imported logic, and no pre-existing arithmetic.
Standard termination methods fail on this system. We document ten distinct proof strategies that succumb to the system’s structural complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, and unchecked recursion. These methods collapse at the same barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats every additive measure ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). This failure is not a defect of the proof search but a necessary feature of any system capable of sequential computation.
To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit. For this fragment, a novel triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and ordinal ranking) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
We state a fundamental conjecture: No relational operator-only TRS can have its full-system termination proved by internally definable methods. We define a “relational” TRS as one capable of internalized sequential computation. This property is shared by Turing machines (via transition functions), lambda calculus (via fixed-point combinators), and recursive functions. All such systems require a recursor to define the “next” step; this recursor inherently creates the duplication barrier.
This result reveals a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), this barrier appears at a more primitive layer: in operator systems possessing only order and recursion, before arithmetic emerges. The system’s inability to prove its own termination from within is a structural limitation of ordered computation itself, paralleling the halting problem but manifesting at the substrate of operator definitions.
The complete Lean 4 formalization ( 7,000 lines) is available at https://github.com/MosesRahnama/OperatorKO7.
1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture regarding the unprovability of termination for relational operator-only systems using internal methods.
Scope. The formal results apply to a guarded safe subrelation. This restriction is necessary because the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ when $\kappa^{M}(a)=0$. The safe fragment employs guarded/context joiners to ensure unique normal forms (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
Contributions
Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. This result supports a certified normalizer with proven totality and soundness. Furthermore, confluence for the safe fragment is established via Newman’s Lemma given verified local confluence. The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
Highlights (formalization summary)
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
  • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
  • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
  • Full Step caveat: A specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails is exhibited, justifying the restriction to the safe subrelation.
  • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
2 Background: TRSs, SN, reachability, and Newman
Standard abstract reduction and term rewriting notions apply [2, 8]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [7], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [1]. Modularity holds in certain linear, non-collapsing combinations [3], while even flat non-linear systems exhibit undecidable reachability and confluence [6]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [4] or (recursive) path orders [5].
3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing.
Rule Head Arity Shape Dup?
R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
Kernel signature (7 constructors)
Terms are generated by the constructors:
$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\,\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
Tiny example (trace consequences)
Using the verified normalizer, we observe:
  • Integrate/delta: $\mathrm{integrate}(\delta\,t)\Rightarrow^{\ast}\mathrm{void}$ (verified).
  • Equality (meta-level consequence under confluence):
    • If $\mathrm{nf}(a)=\mathrm{nf}(b)$ then $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{void}$.
    • Otherwise $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, confluence ensures these outcomes are unique.
Note. Strong normalization and Newman-based confluence are proven for the safe fragment.
4 Strong normalization
The termination argument relies on a triple-lexicographic measure
$\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are formally handled (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). Duplicating branches use a compact MPO-style head precedence.
Step vs SafeStep (measure unification)
For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices. Steps in the full relation are covered by a disjunctive decrease certificate: each step corresponds to either a KO7 lex drop or an MPO triple drop.
Theorem 1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}(t)$.
[…]
DM vs MPO on rules (explicit)
For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$. For eqW, we use a compact MPO-leaning measure where the head precedence orients $\mathrm{eqW}\,a\,b$ strictly above $\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
  • Base-order premise (merge-cancel): in $\mathrm{merge}\,t\,t\to t$, the RHS is strictly smaller than the LHS.
  • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
No-Go for constant bumps on $\kappa$ (generic duplicator)
Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
Let $\kappa$ be a max-depth-style counter and fix any $k\in\mathbb{N}$. For a duplicating rule of the shape $r(S)\to C[S,S]$ (one redex replaced by two occurrences of a subterm $S$), there exists an instance where $\kappa(\text{LHS}){+}k=\kappa(\text{RHS}){+}k$, so no strict lex drop occurs. The same holds for a boolean phase flag alone.
Proof sketch. Choose $S$ with $\kappa(S)\geq 1$ and let $\mathrm{base}$ bound the context. Then $\kappa(\text{LHS})=\mathrm{base}{+}1$ while $\kappa(\text{RHS})=\max(\kappa(C[S]),\kappa(C[S]))=\mathrm{base}{+}1$. Adding a fixed $k$ preserves equality; a single flag does not alter the tie. $\square$
Remark 1.
[…]
Duplication stress identity
For any additive counter $\rho$ that counts a single removed redex and sums subpieces, generic duplicators satisfy
$\rho(\text{after})=\rho(\text{before})-1+\rho(S),$
so there is no strict drop when $\rho(S)\geq 1$. The robust fix uses DM/MPO: replace one element by a multiset of strictly smaller elements (DM), or use RPO/MPO with a precedence/status such that the LHS redex strictly dominates each RHS piece.
Corollary 1 (Strong normalization).
The guarded relation SafeStep for KO7 is strongly normalizing.
Genealogy of failures (why DM/MPO)
Minimal counterpatterns that motivate the multiset/path components are recorded:
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 1).
  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
  • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
The fix is DM/MPO: ensure each RHS piece is strictly smaller than the removed redex in a base order and lift via a multiset/path extension.
5 A certified normalizer
A normalization function is defined by well-founded recursion on $\mu^{3}$. Properties proven in the formalization include:
(Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
(Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
No specific efficiency claim is made: worst-case normalization cost follows the termination witness (cf. the small-term reachability bounds in [1]). Under confluence, decision reduces to one normalization and an equality check.
6 Local confluence and Newman (guarded safe relation)
Local confluence is discharged by joining the finite set of critical pairs. Combining Cor. 1 with Newman’s Lemma [7] yields:
Theorem 2 (Confluence and unique normal forms).
If a relation is strongly normalizing and locally confluent, then it is confluent. Hence every term reduces to a unique normal form.
Instantiation (SafeStep). Combining Cor. 1 (SN for SafeStep) with local-join lemmas yields confluence and unique normal forms for the safe fragment by Newman’s Lemma. In contrast, the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ under $\kappa^{M}(a)=0$; thus full confluence does not hold. The SafeStep relation restricts eqW to cases where arguments are guarded, preventing this divergence.
The star-star join proof follows the standard accessibility (Acc) recursion at the source, with a case split over star shapes (“head step + tail”) and composition via transitivity. The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
Scope and Guarantees (formalization-accurate)
The formal results concern a guarded safe subrelation SafeStep for which we prove:
  • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
Critical-pair coverage
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which is shown to be not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
Table 2: Critical-pair coverage at the root for SafeStep.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta  (unique target void)
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left  (unique target $t$)
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right  (unique target $t$)
$\mathrm{merge}\;t\;t$ localJoin_merge_tt  (unique target $t$)
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero  (unique target $b$)
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ  (unique target $\mathrm{app}\,s\,(\mathrm{rec}\,\Delta\,b\,s\,n)$)
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne  (unique target $\mathrm{integrate}(\mathrm{merge}\,a\,b)$)
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ not locally joinable at root
Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
$\delta$-guard: definition and decidability
Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$. Here $\mathrm{deltaFlag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
[…]
$\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
7 Impossibility results: The Ten Failed Methods
The formal development establishes the failure of ten distinct termination strategies, necessitating the use of the triple-lexicographic measure with Dershowitz-Manna orderings. These failures are not merely heuristic; they are formally witnessed by counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Operational_Incompleteness.lean).
  1. 1.
    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
  2. 2.
    Simple Lexicographic Measures: Pairs like $(\kappa,\mu)$ fail because the primary $\kappa$ component does not strictly decrease on duplicating steps.
  3. 3.
    Polynomial Interpretations: Without external arithmetic coefficients (banned by the operator-only constraint), polynomial measures degrade to additive ones and fail on duplication.
  4. 4.
    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void‘) can toggle the flag from 0 to 1.
  5. 5.
    Pure Ordinal Measures: Shape-blind ordinals fail to distinguish nested $\delta$ structures from their contexts.
  6. 6.
    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like merge_cancel where the result may have equal depth to the input.
  7. 7.
    Naive Multiset Orderings: Standard multiset measures fail without the specific properties of the Dershowitz-Manna order (which allows a single element to be replaced by any number of smaller elements).
  8. 8.
    Local Helper Lemmas: Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
  9. 9.
    Complex Hybrid/Constellation Measures: Attempts to combine measures in ad-hoc ways failed to provide a uniform decrease across all 8 rules.
  10. 10.
    Unchecked Recursion: The raw rec_succ rule itself serves as the ultimate counterexample; without the SafeStep guards, it permits unbounded recursion that no internal measure can bound.
These ten failures collectively demonstrate that the termination barrier is structural: the system’s internal logic cannot capture the necessary invariants without the sophisticated machinery of the triple-lexicographic measure.
Polynomial Impossibility
Consider polynomial interpretations of the form $M(t)\in\mathbb{N}$ for a duplicating rule $f(x)\to g(x,x)$. A polynomial measure assigning $M(f(x))=M(x)+p$ requires $M(x)+p>2M(x)$ for strict decrease, implying $p>M(x)$. This fails for unbounded $M(x)$.
More generally: any polynomial proof for a system with duplicating rules must hardcode specific constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the proof. This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
Internally definable measures
A simple contract for internally defined termination measures structures these negative results:
Definition 1 (Internally definable measure).
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with explicit witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
8 Decidability of Reachability
Theorem 3 (Fixed-target reachability).
In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits. This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
Assumptions (model)
The model assumes a finite first-order TRS over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
Complexity context
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [3], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [6].
9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means
A relational operator-only TRS is one capable of sequential computation: any system that can express “do $X$, then do YY” without external conditions. The terms are equivalent: “relational” = “capable of sequential computation.”
9.1 The Relational Nature of KO7 and Recursors
The classification of KO7 as “relational” is not arbitrary; it describes the system’s capability to internalize sequential control flow. In flat rewrite systems (e.g., group axioms), rules rewrite terms to equivalent terms. In a relational TRS like KO7, the rewrite rules define a computational process.
[…]
$\mathrm{rec}\Delta(b,s,\delta(n))\to\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the transition function $\delta$ of a Turing machine or the fixed-point combinator $Y$ in lambda calculus. Both mechanisms allow a system to execute a sequence of operations of unbounded length.
The Recursor Trap
Any system capable of this sequential computation must have a mechanism to carry the “step” logic ($s$) forward into the next iteration. In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7), this redistribution of $s$ defeats all additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A system that internalizes its own control flow in this way cannot prove its own termination using simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
Escaping the Trap: Meta-Layers vs. Internal Proofs
Standard systems escape this trap by stepping outside the system.
  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally. The system doesn’t prove itself; the compiler proves the system.
  • Turing Machines: We do not ask the tape to prove it halts; we analyze the machine from the outside (and finding that we cannot, we discover the Halting Problem).
KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag). The Conjecture posits that for the full, unguarded system, no such internal proof exists.
The structural barrier
Any operator system capable of ordered computation requires a recursor. A recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 2 (Internally definable measure).
An internally definable termination method for signature $\Sigma$ uses only: simplification orders with fixed precedence/status on $\Sigma$ (LPO/RPO/MPO), DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Conjecture 1 (Full Termination Conjecture for Relational TRSs).
No relational operator-only TRS can have its full-system termination proved by internally definable methods. Specifically: let $R$ be an operator-only TRS with a recursor rule of the form
$\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n)).$
No internally definable measure (Definition 2) proves termination of $R$ when the step argument $s$ is unrestricted. The structural redistribution of $s$ into both the function application and the recursive call defeats every additive, polynomial, or path-order measure that does not import external arithmetic or ordinal axioms. Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
Scope of the conjecture
The conjecture applies to any operator-only TRS capable of sequential computation: any system that can express “do $X$, then do YY” using only rewrite rules, without external axioms, imported arithmetic, or borrowed logic. Systems without such capability (purely ground, non-recursive, or lacking step-iteration) fall outside the conjecture’s scope. The claim is that once a system can compute sequentially without external conditions, its full termination escapes internal proof methods.
Evidence
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta$-phase bit restricts allowable step arguments; without such restriction, the full system admits unbounded recursive depth. The gap between internally definable ranking functions and what would be required cannot be closed without external methods (axioms, encodings, or borrowed arithmetic).
10 Formalization structure
The formal verification is implemented in a proof assistant. The project structure separates the kernel definitions, meta-theory proofs, and impossibility results:
  • Termination proofs: Establishes the per-rule decreases and strong normalization for the safe fragment using the triple-lexicographic measure.
  • Normalization: Defines the normalization function and proves its totality and soundness.
  • Confluence: Implements the Newman engine and the star-star join, relying on local join lemmas.
  • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development (Lean 4,  7,000 LOC; Apache-2.0 licensed) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation. These results yield a certified normalizer and unique normal forms via Newman’s Lemma. An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified. Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.
Appendix A Addendum
A.1 Lean kernel location
The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
  • $\delta$-flag: phase bit dropping on the successor recursion branch.
  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
  • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
Aggregation for $\kappa^{M}$ (union, not sum)
It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
Short witness snippets (toy duplication)
Toy rule: $\mathrm{pair}(\mathrm{s}\,x,\,y)\to\mathrm{pair}(x,\,\mathrm{pair}(y,\,y))$.
DM multiset on sizes. Let $S(x,y)=\mathrm{size}(\mathrm{pair}(\mathrm{s}\,x,\,y))=\mathrm{size}(x)+\mathrm{size}(y)+2$. Then $\mathrm{size}(x)<S(x,y)$ and $\mathrm{size}(y)<S(x,y)$. Use $X=\varnothing$, $Y=\{\mathrm{size}(x),\mathrm{size}(y)\}$, $Z=\{S(x,y)\}$ to conclude $Y<_{\mathrm{DM}}Z$.
MPO triple weight. $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
[…]
Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
A.3 Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
A.4 Module map
The verification logic is distributed across key modules for SN, normalizer, and confluence engine. Local confluence lemmas are proved separately. Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
[…]
Table 3: Per-rule decrease witnesses used in the SafeStep termination proof.
Rule Base component Precedence/Status Witness Source
merge void left $\mu$ (ordinal) (n/a) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) (n/a) Theorem 4.2 Termination
merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
rec succ [^2] This row records the per-step decrease used in the SafeStep proof. It does not assert termination of the unrestricted full relation: rec_succ duplicates the step argument $s$ and is the structural source of the full-system termination barrier (cf. the conjecture/evidence discussion). $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
A.6 Local-join report (Confluence)
Table 4: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
Type Shape Join lemma name Notes
Root merge void void localJoin_merge_void_void both branches to void
Root merge t t localJoin_merge_tt joins at $t$ (context-lifted)
Root integrate (delta t) localJoin_int_delta unique target void
[…]
Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
Root eqW a a, guard ne localJoin_eqW_refl_guard_ne refl blocked; only diff
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne refl blocked
Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
Appendix B Orientation Strategy
This summarizes the exact orientation strategy realized in the proofs:
  • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0).
  • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.