### v4

# Strong Normalization for the Safe Fragment of a Minimal Rewrite Calculus: A Triple-Lexicographic Proof and a Conjecture on Full-Termination Limits for Pure Recursive Calculi (PRC)

Abstract
This paper introduces the concept of a Pure Recursive Calculus (PRC), a minimal computational substrate defined by four strict structural conditions: no variable binding, no external memory, no imported axioms, and an unrestricted step-duplicating recursor. In our PRC framing, this recursor is the mechanism used to represent sequential iteration. We present KO7, a concrete PRC consisting of seven constructors and eight reduction rules, designed to investigate the foundations of self-verifying computation.
Standard termination methods fail on this calculus due to a core structural barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats additive measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth. Furthermore, classical path orders (LPO/MPO) bypass this barrier by relying on the universal Subterm Property, which acts as an imported external axiom. Counting these external path-order methods yields fourteen total strategy families (Section 7). This pattern indicates a structural limitation for calculi of this form rather than a defect of proof search.
To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit. For this fragment, a triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and a computable natural rank) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
We state a central conjecture: For Pure Recursive Calculi (PRC), no internally definable method currently known to us proves full-calculus termination. Unlike Turing machines (which require an external tape), lambda calculus (which requires variable binding), and recursive functions (which import zero and successor as external primitives), a PRC has none of these external scaffolds. On this framing, the termination barrier is positioned at a more primitive layer than those classical settings.
These findings motivate a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), the barrier studied here appears at a more primitive layer: in operator calculi possessing only order and recursion, before arithmetic emerges. In this setting, the difficulty of proving full termination from internal methods parallels halting-style limitations at the substrate of operator definitions.
The complete Lean 4 formalization ($\sim$6,000 LOC) is available at https://github.com/MosesRahnama/OperatorKO7.
1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the unprovability of termination for Pure Recursive Calculi (PRC) using internal methods.
Scope. The formal results apply to a guarded safe subrelation. This restriction is necessary because the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$. The safe fragment employs guarded/context joiners; the Newman local-join hypothesis is explicitly formulated and fully discharged in the artifact (§6).
Why KO7?
[…]
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality.
Contributions
Strong normalization is proven for a guarded fragment of KO7 using a triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and a computable natural rank. This result supports a certified normalizer with proven totality and soundness. Furthermore, a Newman engine for the safe fragment is formalized, and the global local-join hypothesis is fully discharged, yielding unconditional confluence and unique normal forms for the safe fragment.
Furthermore, we apply established proof theory to classify the boundaries of internal termination proving. Drawing on the work of Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi. These methods achieve termination by introducing the universal Subterm Property ($f(t)>t$) as a precondition, which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension). Because these orders import external structural axioms and well-ordering principles, they are outside the internal method class used in this paper. While not a proof of general non-termination, the work includes formally verified impossibility results for fourteen distinct strategy families under duplication. These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
Highlights (formalization summary)
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and a computable natural rank.
  • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
  • Newman (SafeStep): A Newman theorem is formalized for SafeStep; the global local-join hypothesis is discharged in the artifact, yielding unconditional confluence and unique normal forms for the safe fragment.
  • Full Step caveat: A concrete non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is exhibited for the full relation, justifying restriction to a guarded fragment.
  • Impossibility results: Twelve distinct impossibility barriers (including additive, polynomial, weight-based, and precedence-based measures) are formally machine-checked in Impossibility_Lemmas.lean and Conjecture_Boundary.lean, alongside two meta-theoretical barriers.
2 Background: Term Rewriting, SN, reachability, and Newman
Standard abstract reduction and term rewriting notions apply [2, 13]. A term is in normal form if no rule applies. A rewrite calculus is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [11], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating calculi has well-charted complexity: NP-complete for length-reducing calculi (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating calculi [1]. Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear calculi exhibit undecidable reachability and confluence [10]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [5] or (recursive) path orders [7].
3 The KO7 calculus
KO7 is a finite rewrite calculus over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified SafeStep fragment adds guards: R_eq_diff requires $a\neq b$, R_merge_void_left/right require $\delta$-flag tie conditions, R_merge_cancel requires both $\delta$-flag and $\kappa^{M}$ guards, R_rec_zero requires a $\delta$-flag guard on $b$, and R_eq_refl requires $\kappa^{M}(a)=0$. R_merge_cancel is collapsing. (* Note: R_rec_succ duplicates the step argument $s$ (one occurrence on the LHS, two on the RHS). This duplication is the termination stressor: the paper proves strong normalization for the guarded SafeStep fragment, not for the unrestricted full relation.)
Rule Head Arity Shape Dup?
R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
Kernel signature (7 constructors)
Terms are generated by the constructors:
$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\,\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
Tiny example (trace consequences)
Using the verified normalizer, we observe:
  • Integrate/delta: $\mathrm{integrate}(\delta\,t)\Rightarrow^{\ast}\mathrm{void}$ (verified).
  • Strict Syntactic Equality:
    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (provided $\kappa^{M}(a)=0$).
    • If $a\neq b$, then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, arguments inside $\mathrm{eqW}$ are not subject to context closure, enforcing strict syntactic comparison rather than normal-form evaluation.
Note. Strong normalization and Newman-based confluence are proven for the safe fragment.
4 Strong normalization
The canonical termination argument relies on a computable triple-lexicographic measure
$\mu^{3}_{c}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\tau(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna); and (iii) a computable natural payload $\tau$ for non-duplicating ties.
Step vs SafeStep (artifact scope)
For the SafeStep relation, a single unified computable triple lex order $(\delta,\kappa^{M},\tau)$ suffices and is the canonical certificate used in this paper (Meta/ComputableMeasure.lean, with core relation definitions in Meta/SafeStep_Core.lean). For the unrestricted full relation Step, termination is not claimed in this artifact.
Theorem 4.1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}_{c}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}_{c}(t)$.
[…]
DM and $\tau$ roles on rules (explicit)
For rec-zero, the $\kappa^{M}$ component provides a strict DM drop. For merge-cancel, the $\kappa^{M}$ component ties at 0, and the strict decrease is provided by the $\tau$ component. For rec-succ, the outer $\delta\text{-flag}$ drops from $1$ to 0. For eqW-rules in SafeStep (where $\kappa^{M}(a)=0$ is a guard), strict decrease is discharged in the computable $\tau$ component; the more general unguarded lemma in the artifact also includes a $\kappa^{M}(a)\neq 0$ branch discharged by DM.
  • DM premise: in duplicating-sensitive branches, the multiset of RHS pieces is strictly DM-smaller than the removed LHS redex profile.
  • $\tau$ premise: when the $\delta\text{-flag}$ and $\kappa^{M}$ tie, a direct natural-number inequality closes the branch.
No-Go for constant bumps on $\kappa$ (generic duplicator)
Lemma 4.2 (No fixed +k+k or boolean flag orients a generic duplicator).
Let $\kappa$ be a max-depth-style counter and fix any $k\in\mathbb{N}$. For a duplicating rule of the shape $r(S)\to C[S,S]$ (one redex replaced by two occurrences of a subterm $S$), there exists an instance where $\kappa(\text{LHS}){+}k=\kappa(\text{RHS}){+}k$, so no strict lex drop occurs. The same holds for a boolean phase flag alone.
Proof sketch. Choose $S$ with $\kappa(S)\geq 1$ and let $\mathrm{base}$ bound the context. Then $\kappa(\text{LHS})=\mathrm{base}{+}1$ while $\kappa(\text{RHS})=\max(\kappa(C[S]),\kappa(C[S]))=\mathrm{base}{+}1$. Adding a fixed $k$ preserves equality; a single flag does not alter the tie. $\square$
Remark 4.3.
[…]
Duplication stress identity
For any additive counter $\rho$ that counts a single removed redex and sums subpieces, generic duplicators satisfy
$\rho(\text{after})=\rho(\text{before})-1+\rho(S),$
so there is no strict drop when $\rho(S)\geq 1$. The fix used here is the computable triple stack $(\delta\text{-flag},\kappa^{M},\tau)$: DM handles multiset-sensitive branches and $\tau$ resolves ties.
Corollary 4.4 (Strong normalization).
The guarded relation SafeStep for KO7 is strongly normalizing.
Genealogy of failures (why DM/MPO)
Minimal counterpatterns that motivate the multiset/path components are recorded:
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\mathrm{delta}$ structures from their contexts.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 4.2).
  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
  • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
The final artifact path is the computable lex stack $\mu^{3}_{c}=(\delta,\kappa^{M},\tau)$, where DM handles multiset-sensitive branches and $\tau$ resolves tie cases.
5 A certified normalizer
A normalization function is defined by well-founded recursion on SafeStepRev, with well-foundedness certified by the computable measure stack $\mu^{3}_{c}$. The normalizer operates on the root step relation SafeStep (not the context closure SafeStepCtx); a term is in normal form when no root safe rule applies ($\mathrm{NormalFormSafe}(t)\iff\neg\exists u.\;\mathrm{SafeStep}\;t\;u$). Properties proven in the formalization include:
(Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
(Soundness) $\forall t.~\mathrm{SafeStepStar}(t,\mathrm{normalizeSafe}(t))\land\mathrm{NormalFormSafe}(\mathrm{normalizeSafe}(t)).$
No specific efficiency claim is made: worst-case normalization cost follows the termination witness (cf. the small-term reachability bounds in [1]). Under confluence, decision reduces to one normalization and an equality check.
6 Local confluence and Newman (guarded safe relation)
The Newman engine is formalized and the global local-join hypothesis ($\mathsf{locAll}:\forall\,a,\,\mathrm{LocalJoinAt}\;a$) is fully discharged in the artifact. Combined with Cor. 4.4, Newman’s Lemma [11] yields:
Theorem 6.1 (Confluence and unique normal forms).
If a relation is strongly normalizing and locally confluent, then it is confluent. Hence every term reduces to a unique normal form.
Instantiation (SafeStep). The formalization provides a comprehensive library of local-join lemmas over safe root/context shapes, fully discharging the $\mathsf{locAll}$ hypothesis to prove unconditional confluence. In contrast, the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$; thus unrestricted full-step confluence is not claimed. The SafeStep relation introduces guard conditions that avoid this divergence pattern.
The star-star join proof follows the standard accessibility (Acc) recursion at the source, with a case split over star shapes (“head step + tail”) and composition via transitivity. The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
Scope and Guarantees (formalization-accurate)
The formal results concern a guarded safe subrelation SafeStep for which we prove:
  • SN (SafeStep). The computable KO7 triple measure $\mu^{3}_{c}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local-Join Library (SafeStep). Local-join lemmas per root shape and context wrappers are provided; the global hypothesis $\mathsf{locAll}$ is discharged in the artifact (locAll_safe), yielding unconditional confluence and unique NFs.
Critical-pair coverage
The following table records representative safe root configurations with explicit local-join lemmas. We also exhibit a full-step non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ and discharge many guarded eqW cases via context wrappers. Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
Table 2: Critical-pair coverage at the root for SafeStep.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta (unique target void)
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left (unique target $t$)
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right (unique target $t$)
$\mathrm{merge}\;t\;t$ localJoin_merge_tt (unique target $t$)
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero (unique target $b$)
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ (unique target $\mathrm{app}\,s\,(\mathrm{rec}\,\Delta\,b\,s\,n)$)
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne (unique target $\mathrm{integrate}(\mathrm{merge}\,a\,b)$)
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ localJoin_all_safe (eqW/refl branch; target void)
For the $\mathrm{eqW}\,a\,a$ row, the relevant branch inside localJoin_all_safe is discharged via localJoin_of_unique.
$\delta$-guard: definition and decidability
Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$. Here $\mathrm{deltaFlag}:\mathrm{Trace}\to\mathbb{N}$ is a binary root-shape detector for $\mathrm{rec}\Delta\,\_\,\_\,(\mathrm{delta}\,\_)$ defined in the artifact, so the predicate is decidable. Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
[…]
$\mathrm{rec}\Delta\,b\,s\,(\mathrm{delta}\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\Delta\,b\,s\,n)$$\delta\text{-flag}$ drop
7 Impossibility results: The Fourteen Failed Methods
We document the failure of fourteen distinct termination strategy families, motivating the triple-lexicographic measure with Dershowitz-Manna orderings used in the safe fragment. Twelve of these failures are not merely heuristic; they are formally witnessed by machine-checked counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Conjecture_Boundary.lean). The remaining two (local helper lemmas and path orders) are treated as meta-theoretical and methodological barriers.
  1. 1.
    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
  2. 2.
    Simple Lexicographic Measures: Pairs like $(\kappa,\mu)$ fail because the primary $\kappa$ component does not strictly decrease on duplicating steps.
  3. 3.
    Polynomial Interpretations: Without external arithmetic coefficients (banned by the operator-only constraint), polynomial measures degrade to additive ones and fail on duplication.
  4. 4.
    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void’) can toggle the flag from 0 to 1.
  5. 5.
    Pure Ordinal Measures: Shape-blind ordinals fail to distinguish nested $\delta$ structures from their contexts.
  6. 6.
    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like ‘merge_cancel’ where the result may have equal depth to the input.
  7. 7.
    Naive Multiset Orderings: Standard multiset measures fail without the specific properties of the Dershowitz-Manna order (which allows a single element to be replaced by any number of smaller elements).
  8. 8.
    Local Helper Lemmas (Methodological): Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
  9. 9.
    Complex Hybrid/Constellation Measures: Attempts to combine measures in ad-hoc ways failed to provide a uniform decrease across all 8 rules.
  10. 10.
    Unchecked Recursion: The raw ‘rec_succ’ rule itself serves as a counterexample pattern; without the SafeStep guards, it permits unbounded recursion that is not handled by the audited internal families.
  11. 11.
    Path Orders (Meta-Theoretical): Rely on the universal Subterm Property ($f(t)>t$), which injects an external structural axiom and artificial fixed point not present in the pure calculus.
  12. 12.
    Pure Precedence Orders: Measures relying strictly on head constructor precedence fail on collapsing rules like merge_cancel because the RHS can have the exact same head constructor as the LHS.
  13. 13.
    Linear Weight Functions (KBO-style): Purely additive weight functions fail on the duplication in rec_succ regardless of the constants chosen, as the duplicated argument can grow arbitrarily large.
  14. 14.
    Standard Tree Depth: Measuring the maximum depth of the Abstract Syntax Tree (AST) fails because the duplication of $s$ in rec_succ can strictly increase the overall depth of the term.
Taken together, these fourteen failures provide evidence that the termination barrier is structural: within the audited internal families, the necessary invariants are not captured without the triple-lexicographic measure used for SafeStep.
The Boundary Between Code and Meta-Theory
We distinguish between what is machine-checked in the code and what is argued in the meta-theory regarding Path Orders (Item 11). The Lean artifact (Conjecture_Boundary.lean) formally demonstrates that the isolated components of path orders (pure precedence, Item 12, and linear weights, Item 13) mathematically fail to orient the calculus. In classical proofs, successful path-order arguments rely on the universal Subterm Property ($f(t)>t$). While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property. As argued in Section 8.2, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1. The code demonstrates pressure toward such an axiom; the text analyzes its scope.
Polynomial Impossibility
Consider polynomial interpretations of the form $M(t)\in\mathbb{N}$ for a duplicating rule $f(x)\to g(x,x)$. A polynomial measure assigning $M(f(x))=M(x)+p$ requires $M(x)+p>2M(x)$ for strict decrease, implying $p>M(x)$. This fails for unbounded $M(x)$.
More generally for the duplication pattern studied here: polynomial proofs typically require hardcoded constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the orientation argument. This is why polynomial interpretations in this setting are treated as outside the intended internal method class.
8 Decidability of Reachability
Theorem 8.1 (Fixed-target reachability).
In a strongly normalizing and confluent calculus, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits. This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
Assumptions (model)
The model assumes a finite first-order rewrite calculus over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with confluence (hence unique normal forms), this reduces to one normalization and an equality test.
Complexity context
Small-term reachability in terminating calculi ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the known complexity results.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear calculi have undecidable reachability/joinability/confluence [10].
9 Conjecture: Full Termination Limits in Pure Recursive Calculi (PRC)
Sequential Computation
A Pure Recursive Calculus is one capable of sequential computation: any calculus that can express “do $X$, then do YY” without external conditions.
9.1 The Sequential Nature of KO7 and Recursors
The classification of KO7 as a PRC is not arbitrary; it describes the calculus’s capability to internalize sequential control flow. In flat rewrite calculi (e.g., group axioms), rules rewrite terms to equivalent terms. In a sequential calculus like KO7, the rewrite rules define a computational process.
[…]
$\mathrm{rec}\Delta(b,s,\delta(n))\to\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the recursor in Gödel’s System T or the step function of primitive recursive arithmetic. Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
The Recursor Trap
Within this PRC framing, sequential computation requires a mechanism to carry the “step” logic ($s$) into the next iteration. In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7) and formally witnessed in the artifact by the theorem rec_succ_additive_barrier, this redistribution of $s$ defeats additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A calculus that internalizes control flow in this way faces a sharp difficulty when proving full termination with simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
Escaping the Trap: Meta-Layers vs. Internal Proofs
Standard calculi escape this trap by stepping outside the calculus.
  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally. The object calculus does not prove termination by itself; the meta-system checks the required obligations.
  • Turing Machines: We do not ask the tape to prove it halts; we analyze the machine from the outside (and finding that we cannot, we discover the Halting Problem).
KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag and $\kappa^{M}$ side conditions). The conjecture below is scoped to Pure Recursive Calculi (Definition 9.2), not to all operator-only kernels.
The structural barrier
In this PRC setting, ordered computation is represented with a recursor. The recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 9.1 (Internally definable measure).
An internally definable termination method for signature $\Sigma$ uses only: DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Importantly, an internally definable method must be derived strictly from the operational semantics (the rewrite rules). It cannot inject universal structural axioms from the meta-theory. For example, standard simplification orders (LPO/RPO/MPO) are explicitly excluded because they rely on the universal Subterm Property ($f(t)>t$). As demonstrated by Ferreira and Zantema [8], this property acts as an imported precondition axiom, artificially establishing a fixed point (an anchor) that the pure calculus does not possess. Furthermore, as established in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), the well-foundedness of these orders relies on Kruskal’s Tree Theorem, which Gallier [9] notes is a classic independence phenomenon requiring transfinite induction at or above $\varepsilon_{0}$. The distinction is: an internally definable method must be self-contained within the operator language; any proof that borrows structural axioms or well-ordering strength from an external theory is performing meta-level reasoning, not internal reasoning.
Definition 9.2 (Pure Recursive Calculus (PRC)).
A finite first-order rewrite calculus is a Pure Recursive Calculus (PRC) if it satisfies all four of the following conditions:
  1. 1.
    No variable binding. The signature contains no binders; there is no substitution meta-theory and no free variables. All term structure is encoded via constructors, not via variable names.
  2. 2.
    No external memory. The calculus has no external tape, registers, or environment. All computational state is carried within the term itself.
  3. 3.
    No imported axioms. All primitive objects are defined purely as syntactic constructors (e.g., void) governed by the reduction rules, rather than being imported as external mathematical constants with pre-existing axiomatic properties.
  4. 4.
    Unrestricted step-duplicating recursor. The signature contains a recursor rule of the form
    $\mathrm{rec}(b,\,s,\,\sigma(n))\;\to\;f(s,\,\mathrm{rec}(b,\,s,\,n))$
    in which the step argument $s$ appears on both sides and is unrestricted: no guard condition, type annotation, or structural bound is imposed on $s$.
KO7 with full Step is a PRC. Classical Turing machines fail condition 2 (they require an external tape). The untyped lambda calculus fails condition 1 ($\lambda$-abstraction introduces variable binding). Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined). In the sense of this definition, the class targets a more primitive setting than these classical computational models: the termination barrier identified here arises before arithmetic, variable binding, or external memory enter the picture.
9.2 The Subterm Axiom and Artificial Fixed Points
When classical rewriting theory encounters the duplication barrier, it typically deploys simplification orders (such as LPO or MPO). These orders succeed by enforcing the Subterm Property: the universal rule that any term is strictly greater than its subterms ($f(t)>t$).
However, in the context of a Pure Recursive Calculus, this is a meta-theoretical intervention. While the subterm property is a natural consequence of the inductive definition of terms in the meta-theory, importing that inductive structure into the termination measure of a pure calculus constitutes an external intervention. The pure calculus only “knows” what its reduction rules explicitly define. It does not inherently know that $\mathrm{delta}(t)$ is greater than $t$ unless a rule reduces $\mathrm{delta}(t)$ to $t$. By forcing $f(t)>t$ universally across all constructors, simplification orders impose an external structural constraint on the system. They establish an artificial fixed point (an external anchor or “zero”) that the relational logic never defined.
This suggests a structural limitation: proving termination of a pure sequential calculus may require introducing structural axioms at the meta-level. In this analysis, the anchor is introduced from outside the object calculus.
This structural observation aligns with established proof theory. As demonstrated by Ferreira and Zantema [8], the well-foundedness of path orderings depends on assuming the subterm property as a precondition. Furthermore, in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), these orders are formalized as inductively generated relations (least fixed points) whose well-foundedness relies on Kruskal’s Tree Theorem. From the perspective of reverse mathematics, Gallier [9] notes that Kruskal’s theorem is a classic independence phenomenon. As established by Rathjen and Weiermann [12], it requires proof-theoretic strength equivalent to $\Pi^{1}_{1}$-comprehension ($\Pi^{1}_{1}$-CA0) and ordinal bounds reaching into the Veblen hierarchy. Thus, in this framework, orienting a PRC with LPO is treated as meta-level reasoning rather than an internally definable termination proof.
9.3 The Pre-Arithmetic Incompleteness Hierarchy
The conjecture situates KO7 within a three-layer hierarchy of incompleteness results, presented as requiring progressively more external structure:
Table 3: Three layers of computational incompleteness. PRC (Layer 0) requires only an object-level recursor and no external scaffolding; Turing machines (Layer 1) add an external tape and meta-level iteration; Gödel systems (Layer 2) additionally require arithmetic encoding.
Layer Calculus class Minimum requirement Barrier
0 (this work) PRC Order + object-level recursor No known internally definable proof of full termination
1 (Turing) General computation External tape + meta-level transition Cannot decide halting (undecidable)
2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Cannot prove own consistency internally
Why Layer 0 Is Placed Below Layer 1
A classical Turing machine performs sequential computation via a meta-level transition function $\delta:Q\times\Gamma\to Q\times\Gamma\times\{L,R\}$. The iteration (“run for $k$ steps”) is defined outside the machine as a sequence of configuration yields. The tape is external memory; there is no operator within the term language that internalizes the step. In a PRC, by contrast, the recursor $\mathrm{rec}\Delta$ is an object-level constructor: “do the next step” is a rewrite rule applied from inside the term, not a meta-operation on a configuration. This is why a Turing machine fails condition 2 of Definition 9.2 and falls outside the PRC class.
Why Layer 0 Is Placed Below Layer 2
Gödel’s incompleteness theorems require arithmetic sufficient for encoding, at minimum Robinson arithmetic $Q$, with addition, multiplication, zero, and successor as imported primitives. Presburger arithmetic (addition only) and Skolem arithmetic (multiplication only) are both decidable and complete; they escape Gödel because they lack the encoding power of full $Q$. A PRC requires none of this: no imported arithmetic, no Gödel numbering, no self-referential encoding. The termination barrier at Layer 0 arises from structural duplication in the recursor (a constructive phenomenon) rather than from a diagonal self-reference argument. Arithmetic itself is a higher-layer phenomenon: it emerges when the PRC operators are later interpreted (e.g., $\delta$ as successor, $\mathrm{rec}$ as primitive recursion). The Layer 0 barrier exists before that interpretation is imposed. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
Conjecture 9.3 (Scoped Full-Termination Conjecture for Pure Recursive Calculi).
For Pure Recursive Calculi (Definition 9.2), no internally definable measure currently known to us proves full-calculus termination. Specifically: let $R$ be an operator-only calculus with a recursor rule of the form
$\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n)).$
No internally definable measure (Definition 9.1) is currently known to prove termination of such an $R$ when the step argument $s$ is unrestricted. Specifically:
  • Additive and polynomial methods are formally excluded in the audited family. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$; polynomial interpretations without imported arithmetic degrade to additive measures and fail by the same identity (Lemma 4.2). Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
  • The Subterm Axiom Boundary (LPO/RPO/MPO): Standard path orders are explicitly excluded from internally definable methods because they rely on the universal Subterm Property ($f(t)>t$). This property acts as an imported structural axiom, establishing an anchor that the pure calculus does not define via its own rewrite rules. Forcing $f(t)>t$ universally bypasses the duplication obstruction by adding meta-level structure. Under Definition 9.1, importing this axiom violates the “no imported axioms” boundary for PRCs; therefore, such arguments are treated as meta-theoretical rather than internal proofs.
    Note on orientation layers: The canonical SafeStep termination proof in this paper is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean: rec_zero is discharged by DM on $\kappa^{M}$, rec_succ by the $\delta$ drop, and merge_cancel/eqW tie branches by $\tau$. MPO/LPO appear only as external meta-theoretic comparators in the boundary discussion, not as the internal certified certificate.
Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside this scope.
Scope of the conjecture
The conjecture applies to Pure Recursive Calculi (Definition 9.2) under the internal-method boundary above. Calculi that fail any of the four PRC conditions (by introducing external memory, variable binding, imported axioms, or a guarded recursor) may lie outside this scope.
Evidence
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta\text{-flag}$ restricts allowable step arguments; without such restriction, the full calculus admits unbounded recursive depth. At present, no internally definable ranking-function closure is known in this setting without external methods (axioms, encodings, or borrowed arithmetic).
10 Formalization structure
The formal verification is implemented in a proof assistant. The project structure separates the kernel definitions, meta-theory proofs, and impossibility results:
  • Termination proofs: Establishes the per-rule decreases and strong normalization for the safe fragment using the triple-lexicographic measure.
  • Normalization: Defines the normalization function and proves its totality and soundness.
  • Confluence: Implements the Newman engine and the star-star join, relying on local join lemmas.
  • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
All theorem statements currently asserted in the build are formally proven without unproven postulates. Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$). The artifact discharges this parameter via locAll_safe from Meta/Confluence_Safe.lean to obtain unconditional confluentSafe and unique normal forms for SafeStep. The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization for a guarded safe subrelation and formalizing a Newman confluence engine. These results yield a certified normalizer; unique normal forms follow unconditionally for the safe fragment. An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified.
Appendix A Addendum
A.1 Lean kernel location
The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
A.2 Auxiliary orientation sketch (DM/MPO)
This subsection records historical orientation notes used for boundary analysis. It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean. An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
  • $\delta$-flag: phase bit dropping on the successor recursion branch.
  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
  • Tie-break rank $r_{\mathrm{aux}}$: resolves non-duplicating ties when the first two coordinates are equal.
Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
Aggregation for $\kappa^{M}$
It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and $r_{\mathrm{aux}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
Short witness snippets (toy duplication)
Toy rule: $\mathrm{pair}(\mathrm{s}\,x,\,y)\to\mathrm{pair}(x,\,\mathrm{pair}(y,\,y))$.
DM multiset on sizes. Let $S(x,y)=\mathrm{size}(\mathrm{pair}(\mathrm{s}\,x,\,y))=\mathrm{size}(x)+\mathrm{size}(y)+2$. Then $\mathrm{size}(x)<S(x,y)$ and $\mathrm{size}(y)<S(x,y)$. Use $X=\varnothing$, $Y=\{\mathrm{size}(x),\mathrm{size}(y)\}$, $Z=\{S(x,y)\}$ to conclude $Y<_{\mathrm{DM}}Z$.
MPO triple weight. $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
[…]
Figure 1: Auxiliary triple-lex components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $r_{\mathrm{aux}}$.
A.3 Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step has an explicit non-local-join witness at root, eqW void void; accordingly full-step confluence is not claimed.
A.4 Module map
The verification logic is distributed across key modules for SN, normalizer, and confluence engine. Local confluence lemmas are proved separately. Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
A.5 Auxiliary per-rule orientation (historical)
Earlier drafts used auxiliary orientation sketches for boundary exploration. They are historical context only and are not the canonical certified path of this paper.
The canonical per-rule decrease proofs used here are the computable lemmas in Meta/ComputableMeasure.lean (the drop_R_*_c family and measure_decreases_safe_c), with core SafeStep definitions in Meta/SafeStep_Core.lean.
A.6 Local-join report (Confluence)
Table 4: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
Type Shape Join lemma name Notes
Root merge void void localJoin_merge_void_void both branches to void
Root merge t t localJoin_merge_tt joins at $t$ (context-lifted)
Root integrate (delta t) localJoin_int_delta unique target void
[…]
Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
Root eqW a a, guard ne localJoin_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root merge lemma
Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
[…]
Appendix B Orientation Strategy
This summarizes the exact orientation strategy realized in the proofs:
  • All guarded rules are oriented in one computable lex order Lex3c over $(\delta\text{-flag},\kappa^{M},\tau)$.
  • The rule rec_succ is discharged by the outer $\delta\text{-flag}$ drop (1 $\to$ 0).
  • Multiset-sensitive branches (notably rec_zero) are discharged by DM drops on $\kappa^{M}$, while merge_cancel is discharged by $\tau$.
  • Tie branches are discharged by strict inequalities in the computable natural rank $\tau$.
  • Strong normalization follows from well-foundedness of Lex3c via inverse-image and yields wf_SafeStepRev_c.